# Language Bottleneck Models: A Framework for Interpretable Knowledge Tracing and Beyond

### v1

@@ 6 Discussion
Core contributions
Framing Knowledge Tracing as an inverse problem shifts focus from fitting statistical regularities to understanding the underlying reasons for student responses. Language Bottleneck Models operationalize this view by compressing interaction histories into explicit natural-language summaries that maintain high predictive power while being both interpretable and steerable by domain experts.
Why can’t we just prompt GPT-4o directly?
@@
The split encoder-decoder architecture of LBMs offers three key advantages over direct LLM prompting: it creates a global student model with a single latent summary shared across all predictions rather than isolated per-question reasoning; it ensures faithful summaries through a closed-loop decoding objective that penalizes non-predictive summaries; and it provides an explicit interface layer that teachers can read, steer and intervene on.
Wider applicability
@@
LBMs extend beyond education to any task requiring compact, human-readable summaries with predictive power. The minimal ingredients needed are: (1) a sequence of observations about an entity, (2) a need to predict future behaviors of that entity, and (3) value in having interpretable representations. For example, clinical decision support could distill patient data into textual state descriptions that forecast outcomes while remaining auditable; preventive maintenance could compress sensor logs into explanations predicting machine failure; customer success teams could summarize interaction histories to forecast churn.
Limitations
LBMs face several constraints: context length limitations imposed by underlying LLMs, which could be addressed through iterative summarization; the requirement for textual question content rather than just identifiers, limiting applicability to some KT datasets; and the substantial computational cost of running LLMs. We discuss limitations extensively in the Appendix.
Future work
Several promising extensions include: implementing iterative encoding to process longer inputs through sequential chunking; developing active sensing capabilities to select questions that maximize information about students’ knowledge states; adapting LBMs for evolving knowledge states to capture learning and forgetting over time; expanding beyond question-answer inputs to leverage richer data sources; and combining with pedagogical-alignment techniques like LearnLM (team2024learnlm, ). We discuss these points in more details in the Appendix. By bridging the gap between predictive power and human interpretability, Language Bottleneck Models represent a significant step toward AI systems that can genuinely augment human expertise in education and beyond.
Appendix A Introduction to Knowledge Tracing
In this section, we provide a more detailed introduction to Knowledge Tracing. For recent comprehensive reviews of the field, see [1, 25].
Knowledge Tracing (KT) is the task of modeling students’ knowledge states and predicting their performance on future questions based on their historical interactions with educational content. The goal is to dynamically track the acquisition, retention, and application of knowledge as students engage with learning materials.
A.1 Formal Problem Definition
Knowledge Tracing (KT) is fundamentally a sequence modeling problem where the objective is to infer a student’s latent knowledge state based on observed interactions. Formally, let $\mathcal{X}=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{t},a_{t})\}$ denote a sequence of question-answer pairs, where $q_{i}\in\mathcal{Q}$ represents a question from the question space $\mathcal{Q}$, and $a_{i}\in\mathcal{A}$ represents the student’s answer from the answer space $\mathcal{A}$. The goal of KT is to learn a function that predicts the probability of correctly answering a future question $q_{t+1}$ given the historical interaction sequence: $P(a_{t+1}|q_{t+1},\mathcal{X})$.
In practice, each interaction $(q_{i},a_{i})$ in the dataset typically contains richer information. Questions $q_{i}$ often include metadata about the underlying knowledge constructs they assess (e.g., specific skills, concepts, or learning objectives). For example, a mathematics question might be tagged with constructs like "quadratic equations" or "algebraic factoring." Similarly, observed answers $a_{i}$ can vary in granularity—from binary correctness indicators to specific answer choices in multiple-choice questions, or complete textual/code responses for open-ended questions. Many KT datasets also capture additional contextual information such as response time or self-reported confidence scores, which can be leveraged by KT models.
The prediction target can also extend beyond binary correctness. Models may predict specific answer choices a student will select (see for example Task 2 of Eedi’s NeurIPS 2020 Education Challenge [30]), enabling more nuanced understanding of misconceptions. In domains like programming education, models might predict the exact code a student will write [14], capturing finer-grained knowledge states.
A.2 Knowledge Tracing Paradigms
Bayesian Knowledge Tracing (BKT).
Introduced by [5], BKT models student knowledge as a set of binary latent variables corresponding to specific skills or concepts. For each skill, a student is either in a "known" or "unknown" state. BKT employs a Hidden Markov Model to update these probabilities based on observed student responses, accounting for learning, guessing, and slipping in its probabilistic framework. This approach provides interpretable outputs, with each concept having a probability of mastery, but is limited in modeling the complex interdependencies between skills.
Deep Learning Approaches.
Deep Knowledge Tracing (DKT) [20] marked a paradigm shift by using recurrent neural networks to model student knowledge as continuous latent representations. Unlike BKT’s discrete skill probabilities, DKT captures complex patterns and skill relationships through high-dimensional embeddings, significantly improving prediction accuracy. Subsequent models have enhanced this approach, including: Dynamic Key-Value Memory Networks (DKVMN) [34] which incorporate memory components for better concept tracking; Self-Attentive Knowledge Tracing (SAKT) [19] and Attentive Knowledge Tracing (AKT) [6] which leverage attention mechanisms to capture long-range dependencies; and Graph-based Knowledge Tracing models [17, 33] which incorporate structural information into knowledge representations.
A.3 Current Challenges
Interpretability vs. Accuracy Trade-off.
Deep learning approaches typically achieve higher predictive accuracy but at the cost of interpretability. Their latent representations are opaque high-dimensional vectors that provide limited insight into specific concepts a student has mastered or misconceptions they hold. This limits the usability of these models in practice.
Why do we need interpretability?
At first glance, predictive accuracy alone might seem sufficient for KT—an accurate model could, for example, analyze student responses across many questions to identify overall strengths and weaknesses. However, even highly accurate models can yield misleading or fragile predictions if the underlying reasoning is opaque. For instance, [22] showed that predictive accuracy alone might obscure models relying on superficial patterns like repeated questions, limiting practical usability. Thus, without interpretability, accurate models may fail to meaningfully support educators in diagnosing and intervening in students’ learning processes.
Cold-start Problem.
KT systems struggle with new students or questions where historical data is limited or non-existent: for any new question added to the dataset, traditional models require the observation of answers from many students to learn an accurate latent representation for this question. This "cold-start" problem poses significant challenges for the practical deployment of KT systems, particularly in dynamic educational environments where new content and students are regularly introduced.
Appendix B Extended Results
B.1 Encoder/Decoder Variants (Section 5.2.1)
@@ Appendix B Extended Results
Table B1 shows the accuracy of different combinations of encoder-decoder models on the Synthetic dataset. The top row shows the result of using the strongest model evaluated (gpt-4o) as both encoder and decoder. Then, we vary either the encoder or decoder part of the LBM across other LLMs, and report the resulting accuracy. A clear pattern which emerges is that the resulting LBMs are stronger when gpt-4o is used as the encoder instead of the decoder. This implies that the task of accurately capturing knowledge state information in the bottleneck is harder than predicting answers to future questions when provided with a knowledge state summary.
@@
Table B1: Performance of different LBM encoder-decoder combinations on the Synthetic dataset.
@@
B.2 Traditional KT methods on the Eedi dataset
We run the four benchmark KT methods on the Eedi dataset with the same experimental setup as for Table 1 in the main paper (see Section D.2.2 in this appendix for dataset details). Table B2 shows the resulting accuracies, alongside the accuracy found with Direct and LBM models for different LLMs, reproduced from Table 1. The best KT model (AKT) slightly outperform the best LBM and Direct models with gpt-4o, while other KT models reach similar accuracies to LBM and Direct models with gpt models and Gemma3-27B.
Table B2: Average accuracy on the Eedi dataset for different model types. For each student, the accuracy is computed on the last four questions of the trajectory. KT models show mean $\pm$ SEM across 10 runs. Due to computational constraints, Direct and LBM results are reported from single runs.
Model Accuracy
KT Models
DKT $0.647\pm 0.005$
DKVMN $0.669\pm 0.004$
SAKT $0.680\pm 0.007$
AKT $0.720\pm 0.008$
Direct
Qwen2.5-3B 0.347
Qwen2.5-7B 0.647
Gemma3-12B 0.580
Gemma3-27B 0.670
GPT-4o-mini 0.655
GPT-4o 0.703
LBM
Qwen2.5-3B 0.372
Qwen2.5-7B 0.583
Gemma3-12B 0.625
Gemma3-27B 0.650
GPT-4o-mini 0.645
GPT-4o 0.682
B.3 Training LBM Encoders - Difficulty stratification (Section 5.2.3)
To investigate whether the accuracy gains seen during LBM training (cf Figure 6 in the main paper) vary across students in the dataset, we stratify students according to how many misconception they hold: 0, 1, 2 or 3+. Since each misconception represent additional "irregularities" in the student’s response pattern beyond simple mastery of constructs, this effectively stratifies different difficulty levels across students. Table B3 shows the change in accuracy relative to the gpt-4o baseline across difficulty levels.
Table B3: Relative difference in accuracy between trained Gemma3-12B and the GPT-4o baseline, with students grouped by number of misconceptions.
Accuracy (mean ± std) Relative Diff. (%)
0   ($N=22$) 0.98 ± 0.05 0.92 ± 0.12 0.99 ± 0.05 -5.80 0.90 6.70
1   ($N=17$) 0.89 ± 0.09 0.87 ± 0.15 0.91 ± 0.09 -2.00 3.00 5.00
2   ($N=23$) 0.81 ± 0.16 0.80 ± 0.12 0.89 ± 0.08 -0.80 10.20 11.00
3   ($N=135$) 0.77 ± 0.12 0.74 ± 0.14 0.82 ± 0.11 -4.80 5.70 10.50
B.4 Steering LBM behavior
B.4.1 Steering via reward signals
To demonstrate the possibility to steer LBMs’ behavior via the reward signal, we consider an example use-case where a teacher would like the model to pay particular attention to potential misconceptions held by the student. Following the reward-shaping framework of Section 3.3, we augment the training objective with an additional term that explicitly encourages the model to surface misconceptions. Concretely, we set $\Omega(S)=\mathbbm{1}\!\bigl{[}\text{``misconception''}\in\tilde{\mathcal{S}}\bigr{]}$, where $\tilde{\mathcal{S}}$ is the textual bottleneck emitted by the LBM for student state $S$. This binary reward is added to the accuracy term and optimized with GRPO. Figure B1 confirms that the policy quickly internalizes this incentive: after only a handful of training steps, the proportion of bottlenecks that explicitly mention a misconception goes from less than 80% to >95%, demonstrating that the reward function provides an effective lever for shaping higher-level pedagogical behavior.
Figure B1: Presence of the word "misconception" in the LBM’s bottlenecks during training of the encoder model steered towards mentioning student misconceptions via the reward signal.
B.4.2 Augmenting LBMs with student-specific information to complement input data
To demonstrate how an LBM can be actively steered with additional information we run the following ablation on the Synthetic dataset: 1. Each student trajectory originally probes four constructs. For a given trajectory, we remove every question linked to one construct leaving the input data intentionally incomplete. 2. We then run the input data through the encoder and inject a single teacher-supplied sentence describing the student’s mastery of the missing construct directly into the model’s bottleneck representation (e.g., "The student has mastered addition except in the event of misconceptions.).
We repeat this procedure four times—once for each construct—running the LBM both with and without the additional sentence, and report the results Table B4. Naturally, the models complemented with the additional information of the student’s mastery of the construct missing in the input data outperform models provided only with the incomplete data. This small experiment illustrates a key advantage of LBMs: because the LBM compresses evidence into a text-based bottleneck, it can be complemented with additional information absent from the input data.
Table B4: Accuracy of a gpt-4o-based LBM on the Synthetic dataset while removing one construct from the input data, with and without providing information about the missing construct in the bottleneck. Each run is repeated across all four constructs; mean and standard deviation are reported.
0.749 ± 0.007 0.791 ± 0.013
B.4.3 Augmenting with student-specific information to assist LBM training
Finally, we demonstrate how providing additional information to the LBM can assist with the training process. We train two identical LBMs while providing 2 misconceptions either to the encoder as part of the input data, or to the decoder as part of the bottleneck. The models are trained for one epoch on $800$ students. We then evaluate the resulting models without any additional information provided. Table B5 shows the accuracy before/after training for both models. The model where additional information was provided to the encoder reaches $84\%$ accuracy after one training compared to $80\%$ when it is provided to the decoder. This suggest that the additional information facilitates training of the encoder.
Table B5: Relative difference in accuracy between trained Gemma3-12B and the gpt-4o baseline, with students grouped by number of misconceptions.
Before training After 1 epoch training
Information in X 0.802 ± 0.001 0.840 ± 0.004
Information in S 0.794 ± 0.002 0.799 ± 0.006
B.5 Example bottlenecks on the Eedi dataset

### v2

@@ 6 Discussion
Why can’t we just prompt GPT-4o directly? The split encoder-decoder architecture of LBMs offers three key advantages over direct LLM prompting: it creates a global student model with a single latent summary shared across all predictions rather than isolated per-question reasoning; it ensures faithful summaries through a closed-loop decoding objective that penalizes non-predictive summaries; and it provides an explicit interface layer that teachers can read, steer and intervene on.
@@
Wider applicability LBMs extend beyond education to any task requiring compact, human-readable summaries with predictive power. The minimal ingredients needed are: (1) a sequence of observations about an entity, (2) a need to predict future behaviors of that entity, and (3) value in having interpretable representations. For example, clinical decision support could distill patient data into textual state descriptions that forecast outcomes while remaining auditable; preventive maintenance could compress sensor logs into explanations predicting machine failure; customer success teams could summarize interaction histories to forecast churn.
Limitations and Future Work LBMs face several constraints including context length limitations, requirements for textual question content, and substantial computational costs. Future extensions could address these through iterative encoding for longer inputs, active sensing for optimal question selection, adaptation for evolving knowledge states, expansion beyond question-answer data, and integration with pedagogical techniques like LearnLM (Team et al., 2024). These limitations and extensions are discussed in detail in Appendix G.
Appendix
Appendix A Extended Results
A.1 Motivating Observation 2
Figure A1 compares summaries produced by different LLMs when prompted to describe a student’s knowledge state from 50 question-answer pairs from a synthetic dataset (see Section 4 for details). While all models capture high-level skill mastery, only one correctly identifies a misconception (errors with negative numbers) out of the four existing ones, illustrating that standard summarization approaches often miss crucial latent patterns.
Figure A1: Zero-shot knowledge state encoding compared to ground truth. LLM models with different capabilities are prompted to write a summary of the knowledge state (top left panel) of a student, given 50 observed questions and answers provided in text. The ground truth knowledge state (top right panel) describing the student behavior has three main components: constructs mastered, constructs not mastered, and misconceptions. All three models capture the construct mastery correctly, but are not able to capture any misconception, beside o1 which notices the negative numbers misconception (bottom row). Bottom right notations correspond to components in Figure 1.
A.2 Full results
Table A.2 compares the zero-shot performance of LBMs to all the baselines across all datasets.
Note: The surprisingly low accuracy of some LLMs on EEDI and XES3G5M likely reflects the strong class imbalance in these datasets (68% correct in EEDI; 85% in XES3G5M; cf. Table 4) and the fact that both LBMs and Direct models are evaluated zero-shot, preventing them from adjusting to this imbalance. Consequently, models whose predictions skew toward the minority class (“incorrect”) can score below random accuracy. The high AUCs for some of these models indicate that they still capture some meaningful predictive signal.
Table A1: Results across all datasets and methods. We report the mean and standard deviation across N=10 runs for KT and CD models and N=3 runs for LLM-based models (Direct, LBM). XES3G5M-E and XES3G5M-C denote the English and Chinese versions of the XES3G5M dataset, respectively; note that since the KT and CD models do not use the question content, their results are shared across both versions of the dataset. "Zero-shot" methods (LLM Direct; most LBMs) are run zero-shot on a test set; "Trained" methods (KT and CD models, trained LBM) are trained on a train set and evaluated on a test set. For the trained LBM (last row), "E" and "D" stand for Encoder and Decoder respectively. "*" indicates the part of the model that is trained. AUC is unavailable for closed-source LLM models as they require access to the model’s output logits. Results for QIKT on Synthetic and MIRT on EEDI are omitted due to implementation issues.
Synthetic EEDI (Filt.) XES3G5M-E (Filt.) XES3G5M-C (Filt.) Model Type Model Name ACC AUC ACC AUC ACC AUC ACC AUC KT Models (Trained) akt .87±.01 .87±.01 .72±.02 .66±.02 .87±.01 .68±.03 .87±.01 .68±.03 deep irt .82±.02 .82±.01 .68±.02 .61±.02 .85±.01 .64±.03 .85±.01 .64±.03 dkt .82±.02 .81±.02 .66±.02 .57±.02 .85±.01 .64±.03 .85±.01 .64±.03 dkvmn .82±.01 .82±.01 .67±.03 .58±.02 .85±.01 .67±.03 .85±.01 .67±.03 gkt .71±.07 .67±.08 .63±.05 .58±.03 .86±.01 .60±.06 .86±.01 .60±.06 qikt .68±.00 .57±.00 .82±.00 .53±.00 .82±.00 .53±.00 saint .66±.02 .65±.02 .70±.02 .59±.05 .87±.01 .66±.03 .87±.01 .66±.03 sakt .82±.02 .81±.02 .67±.01 .56±.02 .85±.01 .60±.02 .85±.01 .60±.02 simplekt .86±.01 .85±.01 .72±.02 .63±.03 .86±.01 .68±.02 .86±.01 .68±.02 CD Models (Trained) DINA .58±.01 .63±.01 .51±.00 .55±.01 .50±.01 .55±.01 .50±.01 .55±.01 IRT .61±.01 .64±.01 .67±.00 .62±.01 .83±.00 .68±.01 .83±.00 .68±.01 KaNCD .90±.01 .96±.01 .65±.01 .64±.01 .85±.00 .81±.00 .85±.00 .81±.00 MIRT .63±.01 .68±.01 .71±.01 .66±.01 .71±.01 .66±.01 NCDM .90±.01 .96±.00 .66±.01 .67±.00 .85±.00 .80±.00 .85±.00 .80±.00 LLM Direct (Zero-shot) Qwen2.5-3B .56±.00 .79±.00 .35±.00 .54±.00 .19±.00 .56±.00 .19±.00 .55±.00 Qwen2.5-7B .64±.00 .73±.00 .65±.00 .62±.00 .76±.00 .63±.00 .75±.00 .63±.00 gemma-3-12b .63±.00 .96±.00 .58±.00 .71±.00 .30±.00 .71±.00 .29±.00 .70±.00 gemma-3-27b .62±.00 .94±.00 .67±.00 .76±.00 .33±.00 .78±.00 .32±.00 .79±.00 gpt-4o .85±.01 .66±.00 .82±.00 .79±.01 gpt-4o-mini .81±.01 .67±.01 .74±.08 .75±.07 gpt-5 .87±.00 .71±.02 .81±.01 .79±.01 LBM (Zero-shot) Qwen2.5-3B .61±.01 .65±.01 .38±.00 .55±.01 .27±.01 .58±.01 .30±.01 .63±.02 Qwen2.5-7B .65±.01 .69±.01 .58±.01 .55±.01 .75±.01 .63±.00 .74±.01 .62±.02 gemma-3-12b .79±.00 .85±.00 .62±.01 .64±.02 .63±.01 .67±.01 .57±.02 .65±.01 gemma-3-27b .78±.01 .85±.01 .65±.01 .67±.01 .68±.02 .70±.01 .68±.01 .70±.02 gpt-4o .80±.01 .66±.02 .80±.01 .78±.02 gpt-4o-mini .78±.01 .61±.02 .70±.08 .70±.07 gpt-5 .87±.01 .68±.01 .76±.01 .76±.01 LBM (Trained) E: gpt-5, D: gemma-3-12b* .70±.02 .73±.01 .83±.01 .76±.01
A.3 Training LBM Encoders
A.3.1 Difficulty stratification
To investigate whether the accuracy gains seen during LBM training (Figure 5) vary across students in the dataset, we stratify students according to how many misconception they hold: 0, 1, 2 or 3+. Since each misconception represent additional "irregularities" in the student’s response pattern beyond simple mastery of constructs, this effectively stratifies different complexity levels across student knowledge states. Table A2 shows the change in accuracy relative to the GPT-4o baseline across difficulty levels.
Table A2: Relative difference in accuracy between trained Gemma3-12B and the GPT-4o baseline, with students grouped by number of misconceptions.
Accuracy (mean ± std) Relative Diff. (%)
0   ($N=22$) 0.98 ± 0.05 0.92 ± 0.12 0.99 ± 0.05 -5.80 0.90 6.70
1   ($N=17$) 0.89 ± 0.09 0.87 ± 0.15 0.91 ± 0.09 -2.00 3.00 5.00
2   ($N=23$) 0.81 ± 0.16 0.80 ± 0.12 0.89 ± 0.08 -0.80 10.20 11.00
3   ($N=135$) 0.77 ± 0.12 0.74 ± 0.14 0.82 ± 0.11 -4.80 5.70 10.50
The model shows $10-11\%$ increased accuracy relative to GPT-4o for students with 2 or 3+ misconceptions, compared to just $5-7\%$ for students with 0-1 misconceptions. This suggests the encoder becomes particularly better at handling complex knowledge states.
A.3.2 Evolution of misconception detection and false positive
Figure A2 shows the evolution of the misconception detection rate and misconception false positive counts over the training of the encoder from §4.2.
Figure A2: Evolution of misconception detection and false positive over encoder training.
A.4 Steering LBM behavior
A.4.1 Augmenting with student-specific information to assist LBM training
Finally, we demonstrate how providing additional information to the LBM can assist with the training process. On the Synthetic dataset, we train two identical LBMs (Gemma-12B trainable encoder, Gemma-27B frozen decoder) while providing 2 misconceptions either to the encoder as part of the input data, or to the decoder as part of the bottleneck. The models are trained for one epoch on $800$ students. We then evaluate the resulting models without any additional information provided. Table A3 shows the accuracy before/after training for both models. The model where additional information was provided to the encoder during training reaches $84\%$ accuracy after training compared to $80\%$ when it is provided to the decoder. This suggest that the additional information facilitates training of the encoder.
Table A3: Result of training on the Synthetic dataset when providing misconception information about each student during training, either in the encoder input or in the bottleneck.
Before training After 1 epoch training
Information in X 0.802 ± 0.001 0.840 ± 0.004
Information in S 0.794 ± 0.002 0.799 ± 0.006
A.4.2 Steering via reward signals
To demonstrate the possibility to steer LBMs’ behavior via the reward signal, we consider an example use-case where a teacher would like the model to pay particular attention to potential misconceptions held by the student. Following the reward-shaping framework of Section 3.3, we augment the training objective with an additional term that explicitly encourages the model to surface misconceptions. Concretely, we set $\Omega(S)=\mathbbm{1}\!\bigl[\text{``misconception''}\in\tilde{\mathcal{S}}\bigr$], where $\tilde{\mathcal{S}}$ is the textual bottleneck emitted by the LBM for student state $S$. This binary reward is added to the accuracy term and optimized with GRPO. Figure A3 confirms that the policy quickly internalizes this incentive: after only a handful of training steps, the proportion of summaries that explicitly mention a misconception goes from less than 80% to >95%, demonstrating that the reward function provides an effective lever for shaping higher-level pedagogical behavior.
Figure A3: Presence of the word "misconception" in the LBM’s summaries during training of the encoder model steered towards mentioning student misconceptions via the reward signal.
A.4.3 Augmenting LBMs with student-specific information to complement input data
To demonstrate how an LBM can be actively steered with additional information we run the following ablation on the Synthetic dataset: 1. Each student trajectory originally probes four constructs. For a given trajectory, we remove every question linked to one construct leaving the input data intentionally incomplete. 2. We then run the input data through the encoder and inject a single teacher-supplied sentence describing the student’s mastery of the missing construct directly into the model’s bottleneck representation (e.g., "The student has mastered addition except in the event of misconceptions").
We repeat this procedure four times—once for each construct—running the LBM both with and without the additional sentence, and report the results Table A4. Naturally, the models complemented with the additional information of the student’s mastery of the construct missing in the input data outperform models provided only with the incomplete data. This small experiment illustrates a key advantage of LBMs: because the LBM compresses evidence into a text-based summary, it can be complemented with additional information absent from the input data.
Table A4: Accuracy of a gpt-4o-based LBM on the Synthetic dataset while removing one construct from the input data, with and without providing information about the missing construct in the bottleneck. Each run is repeated across all four constructs; mean and standard deviation are reported.
0.749 ± 0.007 0.791 ± 0.013
A.5 Ablation Experiments
A.5.1 Encoder/Decoder Variants
@@
Table A5 shows the accuracy of different combinations of encoder-decoder models on the Synthetic dataset. The top row shows the result of using the strongest model evaluated (gpt-4o) as both encoder and decoder. Then, we vary either the encoder or decoder part of the LBM across other LLMs, and report the resulting accuracy. A clear pattern which emerges is that the resulting LBMs are stronger when gpt-4o is used as the encoder instead of the decoder. This implies that the task of accurately capturing knowledge state information in the bottleneck is harder than predicting answers to future questions when provided with a knowledge state summary.
@@
Table A5: Performance of different LBM encoder-decoder combinations on the Synthetic dataset.
@@
A.5.2 Importance of Knowledge Concepts in the input
Table LABEL:app:tab:ablation compares the accuracy of different LBMs with vs without knowledge conception information in the input prompt. The performance generally remains similar to the KC version, and it even increases for most models on the EEDI dataset. This demonstrates that LBMs do not fundamentally require KC information.
Table A6: Comparison of LBM performance with vs without knowledge concept (KC) information in the input. Bold indicates a difference in accuracy no more than $3\%$ below the full input.
Synthetic EEDI (Filtered)
w/ KC w/o KC $\Delta$ w/ KC w/o KC $\Delta$
Qwen2.5-3B .61 ± .01 .62 ± .05 +.00 .38 ± .00 .38 ± .02 +.01
Qwen2.5-7B .65 ± .01 .66 ± .02 +.01 .58 ± .01 .58 ± .01 +.00
gemma-3-12b .79 ± .00 .73 ± .00 -.06 .62 ± .01 .60 ± .02 -.02
gemma-3-27b .78 ± .01 .75 ± .01 -.03 .65 ± .01 .69 ± .03 +.03
gpt-4o-mini .78 ± .01 .76 ± .00 -.01 .61 ± .02 .65 ± .03 +.04
gpt-4o .80 ± .01 .76 ± .01 -.04 .66 ± .02 .68 ± .01 +.02
A.6 Example bottlenecks on the Eedi dataset